# Tracing Nitrogen Enrichment across Cosmic Time with JWST

### v2

@@
11institutetext: Università di Firenze, Dipartimento di Fisica e Astronomia, via G. Sansone 1, 50019 Sesto Fiorentino, Florence, Italy 22institutetext: INAF – Arcetri Astrophysical Observatory, Largo E. Fermi 5, I-50125, Florence, Italy 33institutetext: European Southern Observatory, Karl-Schwarzschild Straße 2, D-85748 Garching bei München, Germany 44institutetext: School of Astronomy, Institute for research in fundamental sciences (IPM), Tehran, P.O. Box 19395-5531, Iran 55institutetext: Department of Physics, University of Arkansas, 226 Physics Building, 825 West Dickson Street, Fayetteville, AR 72701, USA 66institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, I-56126 Pisa, Italy 77institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh, EH9 3HJ, UK 88institutetext: Centre for Astrophysics Research, Department of Physics, Astronomy and Mathematics, University of Hertfordshire, Hatfield AL10 9AB, UK 99institutetext: AURA for European Space Agency, Space Telescope Science Institute, 3700 San Martin Drive. Baltimore, MD, 21210 1010institutetext: Cavendish Laboratory, University of Cambridge, 19 $J$ Thomson Avenue, Cambridge, CB3 0HE, UK 1111institutetext: Kavli Institute for Cosmology, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 1212institutetext: University of Trento, Via Sommarive 14, I-38123 Trento, Italy
@@
We present a comprehensive analysis of the nitrogen-to-oxygen (N/O) abundance ratio in sample of $\sim 500$ star-forming galaxies at redshift $z\sim 1-6$, with a median redshift of  ¡ $z$ ¿ = 2.7, using deep JWST/NIRSpec spectroscopy. Leveraging detections of faint auroral emission lines in 76 galaxies at $z>1$ from both the MARTA survey and a large compilation of high-redshift literature objects, we derive direct electron temperature-based abundances for nitrogen and oxygen using rest-frame optical lines. We establish the first high-redshift calibrations of strong-line N/O diagnostics based on ‘direct’ abundance measurements, finding no significant evolution for both the N2O2 = $\log$([N ii] $\lambda$6585/[O ii] $\lambda\lambda$3727,3729) and N2S2 = log([N ii] $\lambda$6585/[S ii] $\lambda\lambda$6717,6731) diagnostics compared to local realisations. We then investigate the N/O–O/H relation across cosmic time using both ‘direct’ abundances and strong-lines based measurements (additional 430 galaxies). We find evidence for mild but systematic nitrogen enhancement at high redshift: galaxies at $z>1$ exhibit N/O ratios elevated by $\sim$0.18 dex (median offset) at fixed O/H compared to local trends, with a more pronounced enhancement at low metallicity (i.e. 12 + log(O/H) $\lesssim$ 8.1), where the offset reaches up to $\sim$0.3-0.4 dex. We consider several scenarios to explain the observed trends, including bursty star formation, differential metal loading, and inflows of pristine gas. Our results provide the most extensive confirmation of elevated N/O ratios at high-redshift to date based on rest-optical diagnostics and within a self-consistent frame.
@@ 1 Introduction
Beyond absolute abundances, the relative abundance patterns of individual elements encode information on the timescales and sources of chemical enrichment, and are therefore powerful diagnostics of the physical processes driving galaxy evolution. Among these, the nitrogen-to-oxygen ratio (N/O) plays a particularly important role. In fact, oxygen is an $\alpha$-element, produced predominantly through helium fusion in short-lived massive stars ($M\gtrsim 8,M_{\odot}$), which end their lives as core-collapse supernovae, enriching the ISM with oxygen. Nitrogen, on the other hand, can be produced through two main channels, commonly referred to as primary and secondary production. In the primary channel, nitrogen originates from carbon and oxygen freshly synthesized within the star itself: these newly formed nuclei are mixed into the hydrogen-burning regions and converted into nitrogen through the CNO cycle. As a result, the nitrogen yield in this case is largely independent of the star’s initial metallicity. Primary nitrogen is thought to be predominantly produced by massive stars on short timescales (tens of Myr), where rotational and convective mixing processes are efficient (e.g., Henry et al., 2000; Meynet & Maeder, 2002; Chiappini et al., 2006). In contrast, the secondary channel relies on carbon and oxygen already present in the progenitor gas cloud. In this case, the CNO cycle transforms pre-existing heavy elements into nitrogen, leading to a metallicity-dependent yield (e.g., Henry et al., 2000; Vincenzo et al., 2016; Maiolino & Mannucci, 2019). This contribution arises mainly from low- and intermediate-mass stars (LIMS hereafter; $M\lesssim 8M_{\odot}$) during their asymptotic giant branch (AGB) phase, on much longer enrichment timescales (hundreds of Myr to several Gyr; Nomoto et al. 2013; Ventura et al. 2013).
As a consequence, the N/O abundance ratio is a sensitive probe of the relative importance of rapid, primary production in massive stars versus delayed, secondary production in AGB stars, and therefore provides crucial insights into the timing and efficiency of chemical enrichment processes driving galaxy evolution (e.g., Vangioni et al., 2018; Vincenzo & Kobayashi, 2018). In fact, the detailed shape and normalization of the N/O-O/H relation, as well as its scatter, are influenced by several factors, including the star-formation history, the initial mass function, metallicity-dependent stellar yields, gas inflows and outflows, in addition to the time delay between oxygen enrichment from massive stars and nitrogen release from LIMS (e.g., Tinsley, 1980; Edmunds, 1990; Maeder, 1992; Meynet & Maeder, 2002; Ventura et al., 2013; Vincenzo et al., 2016; Henry et al., 2000; Arellano-Córdova et al., 2025a).
@@
Direct determination of nitrogen and oxygen abundances at high redshifts have long been limited by observational challenges: the rest-frame optical lines that enable such measurements are shifted into the near-infrared at $z\gtrsim 2$, and the required auroral lines are intrinsically faint. Analyses based on strong-line ratios that correlate with N/O, such as N2O2 - defined as $\log(\text{[{N}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda 6584/\text{[{O}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda\lambda 3727,3729)$ - and N2S2 - defined as $\log(\text{[{N}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda 6584/\text{[{S}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda\lambda 6717,6731)$ - generally find elevated N/O at fixed metallicity relative to $z\sim 0$ samples (e.g., Masters et al., 2014; Strom et al., 2017; Hayden-Pawson et al., 2022) The fundamental limitation of these efforts is that strong-line diagnostics are inherently sensitive to additional parameters, especially the ionisation parameter and abundance ratios like S/O (e.g., Dopita et al., 2016; Pérez-Montero et al., 2021; Scholte et al., 2025). Without detections of faint auroral lines and corresponding electron temperature estimates, it has remained unclear whether observed trends reflect genuine chemical differences or evolving ISM conditions.
@@
The sensitivity and wavelength coverage of JWST/NIRSpec (Jakobsen et al., 2022) now allow robust abundance determinations based on auroral-line detections in galaxies out to $z\gtrsim 10$ (e.g., Bunker et al., 2023; Cameron et al., 2023; Castellano et al., 2024; Marques-Chaves et al., 2024; Napolitano et al., 2024; Navarro-Carrera et al., 2024; Schaerer et al., 2024; Topping et al., 2024a; Curti et al., 2025b).
@@
At intermediate redshift ($z\sim 2$), these new observational capabilities open the possibility of reassessing the N/O–O/H relation with unprecedented precision (Sanders et al., 2023; Strom et al., 2023; Rogers et al., 2024; Welch et al., 2024; Arellano-Córdova et al., 2025b; Curti et al., 2025a; Rogers et al., 2025; Sanders et al., 2025; Scholte et al., 2025; Topping et al., 2025a; Welch et al., 2025).
@@
At even earlier epochs, JWST spectroscopy has recently uncovered surprising evidence for extreme nitrogen enrichment, indicative of super-solar N/O ratios at sub-solar O/H, well above the expectations from standard chemical evolution and nucleosynthetic timescales (Bunker et al., 2023; Cameron et al., 2023; Isobe et al., 2023; Jones et al., 2023; Castellano et al., 2024; Ji et al., 2024; Navarro-Carrera et al., 2024; Schaerer et al., 2024; Topping et al., 2024b; Curti et al., 2025b; Ji et al., 2025; Napolitano et al., 2025; Topping et al., 2025b; Stiavelli et al., 2025; Zhang et al., 2025). Standard chemical evolution models cannot fully account for this nitrogen enhancement. Proposed alternative channels include enrichment by fast-rotating low-metallicity stars sampled from a top-heavy IMF, very massive stars (VMS; $M_{*}>100M_{\odot}$), Wolf-Rayet stars, early AGB enrichment combined with gas flows, bursty star formation history, pristine gas inflows, differential outflows, and rare events such as tidal disruptions (Vink, 2023; Charbonnel et al., 2023; D’Antona et al., 2023; Nagele & Umeda, 2023; Watanabe et al., 2024; Rizzuti et al., 2024; Arellano-Córdova et al., 2025a; Bhattacharya & Kobayashi, 2025; McClymont et al., 2025; Rizzuti et al., 2025).
@@
However, it is important to note that both these individual and stacked measurements of nitrogen enhancements obtained so far with JWST have been derived from rest-frame UV line ratios, such as N iii] $\lambda\lambda$1747,1749, N iv] $\lambda$1486, and O iii] $\lambda$1666. These lines generally probe different, denser, and higher ionisation regions from those probed by rest-frame optical ratios. This effect may introduce systematic differences when compared to optical abundance diagnostics (e.g. Acharyya et al., 2019; Byler et al., 2020; Mingozzi et al., 2022; Martinez et al., 2025). Moreover, pre-JWST studies could be subject to strong selection biases, as ground-based detections of nitrogen lines at high $z$ require extremely bright, massive, and luminous galaxies. This observational bias favours systems with intrinsically bright nitrogen line emission, making it difficult to assess whether such N/O enhancement is representative of the general galaxy population. By contrast, recent JWST measurements using rest-frame optical nitrogen lines at $z\lesssim 4.5$ find either no significant deviation from the local and lower-redshift N/O–O/H relations (Arellano-Córdova et al., 2025a; Rogers et al., 2025; Stiavelli et al., 2025) or mild enhancement (Arellano-Córdova et al., 2025b), except in the case of clearly exceptional systems such as the Sunburst Arc (Welch et al., 2025).
@@ 2 Data
We aim to select the largest possible sample of high-redshift galaxies for which reliable abundance can be measured, either with the direct method or using strong line calibrations tailored to the high-redshift ISM conditions. For comparison, we also assembled an extensive compilation of local galaxies and H ii region spectra, selected according to the same criteria. In particular, in Section 2.1 we describe the selection of galaxies with auroral-line detections, including both high-redshift objects observed with JWST, as well as local galaxies with auroral-line detections enabling direct metallicity measurements. In Section 2.2, we discuss the selection of galaxies where we use strong-line abundance diagnostics, comprising several ground-based spectroscopic surveys targeting Cosmic Noon, a large compilation of archival data observed with JWST, and local galaxies from large spectroscopic surveys. The redshift distribution of the high-redshift sample is shown in Figure 1, demonstrating that the bulk of our objects lie between $z\sim 2$ and $z\sim 4$ (median redshift $z=2.7$).
@@ 2.1.1 The MARTA survey
Details on the spectral properties, fitting strategy, and validation are provided in C25. We required all galaxies to have robust (S/N ¿ 3) detections of the optical [N ii] $\lambda$6585 and [O ii] $\lambda\lambda$3727,3729 emission lines and imposed S/N cuts on the Balmer lines used for the dust correction (as described in Section 3.1). These criteria yield a total of 27 galaxies meeting all requirements. Among these, ten MARTA galaxies show robust (S/N ¿ 3) detections of one or more auroral lines, allowing for direct determinations of the oxygen and nitrogen abundances.
Figure 1: Redshift distribution of the 506 galaxies for which we measure O/H and N/O in the high-$z$ sample. JWST galaxies with auroral-line detections are shown in light-blue; JWST objects without auroral-line detections (strong-line measurements only) are displayed in teal; objects from ground-based surveys (strong-line measurements from MOSDEF and KLEVER) are plotted in green.

### v3

@@
11institutetext: Università di Firenze, Dipartimento di Fisica e Astronomia, via G. Sansone 1, 50019 Sesto Fiorentino, Florence, Italy 22institutetext: INAF – Arcetri Astrophysical Observatory, Largo E. Fermi 5, I-50125, Florence, Italy 33institutetext: European Southern Observatory, Karl-Schwarzschild Straße 2, D-85748 Garching bei München, Germany 44institutetext: INAF – Bologna Astrophysical Observatory, Via Piero Gobetti, 93/3, 40129, Bologna, Italy 55institutetext: School of Astronomy, Institute for research in fundamental sciences (IPM), Tehran, P.O. Box 19395-5531, Iran 66institutetext: Department of Physics, University of Arkansas, 226 Physics Building, 825 West Dickson Street, Fayetteville, AR 72701, USA 77institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, I-56126 Pisa, Italy 88institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh, EH9 3HJ, UK 99institutetext: Centre for Astrophysics Research, Department of Physics, Astronomy and Mathematics, University of Hertfordshire, Hatfield AL10 9AB, UK 1010institutetext: AURA for European Space Agency, Space Telescope Science Institute, 3700 San Martin Drive. Baltimore, MD, 21210 1111institutetext: Cavendish Laboratory, University of Cambridge, 19 $J$ Thomson Avenue, Cambridge, CB3 0HE, UK 1212institutetext: Kavli Institute for Cosmology, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 1313institutetext: University of Trento, Via Sommarive 14, I-38123 Trento, Italy
@@
We present a comprehensive analysis of the nitrogen-to-oxygen (N/O) abundance ratio in a sample of $\sim 660$ star-forming galaxies at redshift $z\sim 1-6$, with a median redshift of $\langle z\rangle$ = 3.0, using deep JWST/NIRSpec spectroscopy. Leveraging detections of faint auroral emission lines in 92 galaxies at $z>1$ from both the MARTA survey and a large compilation of high-redshift literature objects, we derive direct electron temperature-based abundances for nitrogen and oxygen using rest-frame optical lines. We establish the first high-redshift calibrations of strong-line N/O diagnostics based on ‘direct’ abundance measurements, finding a mild evolution for the N2O2 = $\log$([N ii] $\lambda$6585/[O ii] $\lambda\lambda$3727,3729) diagnostic and no clear evolution for the N2S2 = log([N ii] $\lambda$6585/[S ii] $\lambda\lambda$6717,6731) diagnostic compared to local realisations. We then investigate the N/O–O/H relation across cosmic time using both ‘direct’ abundances and strong-lines based measurements (additional 535 galaxies). We find evidence for mild but systematic nitrogen enhancement at high redshift: galaxies at $z>1$ exhibit N/O ratios elevated by $\sim$0.18 dex (median offset) at fixed O/H compared to local trends, with a more pronounced enhancement at low metallicity (i.e. 12 + log(O/H) $\lesssim$ 8.1), where the offset reaches up to $\sim$0.4-0.5 dex. We consider several scenarios to explain the observed trends, including bursty star formation, differential metal loading, and inflows of pristine gas. Our results provide the most extensive confirmation of elevated N/O ratios at high-redshift to date based on rest-optical diagnostics and within a self-consistent frame.
@@ 1 Introduction
Beyond absolute abundances, the relative abundance patterns of individual elements encode information on the timescales and sources of chemical enrichment, and are therefore powerful diagnostics of the physical processes driving galaxy evolution. Among these, the nitrogen-to-oxygen ratio (N/O) plays a particularly important role. In fact, oxygen is an $\alpha$-element, produced predominantly through helium fusion in short-lived massive stars ($M\gtrsim 8,M_{\odot}$), which end their lives as core-collapse supernovae, enriching the ISM with oxygen. Nitrogen, on the other hand, can be produced through two main channels, commonly referred to as primary and secondary production. In the primary channel, nitrogen originates from carbon and oxygen freshly synthesized within the star itself: these newly formed nuclei are mixed into the hydrogen-burning regions and converted into nitrogen through the CNO cycle. As a result, the nitrogen yield in this case is largely independent of the star’s initial metallicity. Primary nitrogen can be produced both by massive stars on short timescales–tens of Myr–where fast rotational and convective mixing processes are highly efficient (e.g., Henry et al., 2000; Meynet & Maeder, 2002; Chiappini et al., 2006), and by low- and intermediate-mass stars (LIMS hereafter; $M\lesssim 8M_{\odot}$), during their asymptotic giant branch (AGB) phase (e.g., Vincenzo et al., 2016). Such production channel happens on much longer enrichment timescales–hundreds of Myr to several Gyr. The relative importance of each channel depends on the star formation history, IMF, and metallicity evolution of the galaxy (e.g., Romano et al., 2019). In contrast, the secondary channel relies on carbon and oxygen already present in the progenitor gas cloud. In this case, the CNO cycle transforms pre-existing heavy elements into nitrogen, leading to a metallicity-dependent yield (e.g., Henry et al., 2000; Vincenzo et al., 2016; Maiolino & Mannucci, 2019). This contribution arises mainly from LIMS in their AGB phase (Kobayashi et al., 2011; Nomoto et al., 2013; Ventura et al., 2013).
As a consequence, the N/O abundance ratio provides crucial insights into the timing and efficiency of chemical enrichment processes driving galaxy evolution (e.g., Vangioni et al., 2018; Vincenzo & Kobayashi, 2018). In fact, the detailed shape and normalization of the N/O-O/H relation, as well as its scatter, are influenced by several factors, including the star-formation history, the initial mass function, metallicity-dependent stellar yields, gas inflows and outflows, in addition to the time delay between oxygen enrichment from massive stars and nitrogen release from LIMS (e.g., Tinsley, 1980; Edmunds, 1990; Maeder, 1992; Meynet & Maeder, 2002; Ventura et al., 2013; Vincenzo et al., 2016; Henry et al., 2000; Arellano-Córdova et al., 2025a).
@@
Direct determination of nitrogen and oxygen abundances at high redshifts have long been limited by observational challenges: the rest-frame optical lines that enable such measurements are shifted into the near-infrared at $z\gtrsim 2$, and the required auroral lines are intrinsically faint.
@@
Analyses based on strong-line ratios that correlate with N/O, such as N2O2 - defined as $\log(\text{[{N}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda 6584/\text{[{O}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda\lambda 3727,3729)$ - and N2S2 - defined as $\log(\text{[{N}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda 6584/\text{[{S}\,{\sc{ii}}]{}{}{}{}}\penalty 10000\ \lambda\lambda 6717,6731)$ - generally find elevated N/O at fixed metallicity relative to $z\sim 0$ samples (e.g., Masters et al., 2014; Strom et al., 2017; Hayden-Pawson et al., 2022) The fundamental limitation of these efforts is that strong-line diagnostics are inherently sensitive to additional parameters, especially the ionisation parameter and abundance ratios like S/O (e.g., Dopita et al., 2016; Pérez-Montero et al., 2021; Scholte et al., 2025). Without detections of faint auroral lines and corresponding electron temperature estimates, it has remained unclear whether observed trends reflect genuine chemical differences or evolving ISM conditions.
@@
The sensitivity and wavelength coverage of JWST/NIRSpec (Jakobsen et al., 2022) now allow robust abundance determinations based on auroral-line detections in galaxies out to $z\gtrsim 10$ (e.g., Bunker et al., 2023; Cameron et al., 2023; Castellano et al., 2024; Marques-Chaves et al., 2024; Schaerer et al., 2024; Topping et al., 2024; Curti et al., 2025b; Napolitano et al., 2025; Navarro-Carrera et al., 2025).
@@
At intermediate redshift ($z\sim 2$), these new observational capabilities open the possibility of reassessing the N/O–O/H relation with unprecedented precision (Sanders et al., 2023; Strom et al., 2023; Rogers et al., 2024; Welch et al., 2024; Arellano-Córdova et al., 2025b; Curti et al., 2025a; Sanders et al., 2025; Scholte et al., 2025; Topping et al., 2025a; Welch et al., 2025; Rogers et al., 2026).
@@
At even earlier epochs, JWST spectroscopy has recently uncovered surprising evidence for extreme nitrogen enrichment, indicative of super-solar N/O ratios at sub-solar O/H, well above the expectations from standard chemical evolution and nucleosynthetic timescales (Bunker et al., 2023; Cameron et al., 2023; Isobe et al., 2023; Jones et al., 2023; Castellano et al., 2024; Ji et al., 2024; Schaerer et al., 2024; Topping et al., 2024; Curti et al., 2025b; Napolitano et al., 2025; Navarro-Carrera et al., 2025; Topping et al., 2025b; Stiavelli et al., 2025; Ji et al., 2026; Zhang et al., 2026). Standard chemical evolution models cannot fully account for this nitrogen enhancement. Proposed alternative channels include enrichment by fast-rotating low-metallicity stars sampled from a top-heavy IMF, very massive stars (VMS; $M_{*}>100M_{\odot}$), Wolf-Rayet stars, early AGB enrichment combined with gas flows, bursty star formation history, pristine gas inflows, differential outflows, and rare events such as tidal disruptions (Vink, 2023; Charbonnel et al., 2023; D’Antona et al., 2023; Nagele & Umeda, 2023; Watanabe et al., 2024; Rizzuti et al., 2024; Arellano-Córdova et al., 2025a; Rizzuti et al., 2025; Bhattacharya & Kobayashi, 2026; McClymont et al., 2026).
@@
However, it is important to note that both these individual and stacked measurements of nitrogen enhancements obtained so far with JWST have been derived from rest-frame UV line ratios, such as N iii] $\lambda\lambda$1747,1749, N iv] $\lambda$1486, and O iii] $\lambda$1666. These lines generally probe different, denser, and higher ionisation regions from those probed by rest-frame optical ratios. This effect may introduce systematic differences when compared to optical abundance diagnostics (e.g. Acharyya et al., 2019; Byler et al., 2020; Mingozzi et al., 2022; Martinez et al., 2025). Moreover, pre-JWST studies could be subject to strong selection biases, as ground-based detections of nitrogen lines at high $z$ require extremely bright, massive, and luminous galaxies. This observational bias favours systems with intrinsically bright nitrogen line emission, making it difficult to assess whether such N/O enhancement is representative of the general galaxy population. By contrast, recent JWST measurements using rest-frame optical nitrogen lines at $z\lesssim 4.5$ find either no significant deviation from the local and lower-redshift N/O–O/H relations (Arellano-Córdova et al., 2025a; Stiavelli et al., 2025; Rogers et al., 2026) or mild enhancement (Arellano-Córdova et al., 2025b), except in the case of clearly exceptional systems such as the Sunburst Arc (Welch et al., 2025).
@@ 2 Data
We aim to select the largest possible sample of high-redshift galaxies for which reliable abundance can be measured, either with the direct method or using strong line calibrations tailored to the high-redshift ISM conditions. For comparison, we also assembled an extensive compilation of local galaxies and H ii region spectra, selected according to the same criteria. In particular, in Section 2.1 we describe the selection of galaxies with auroral-line detections, including both high-redshift objects observed with JWST, as well as local galaxies with auroral-line detections enabling direct metallicity measurements. In Section 2.2, we discuss the selection of galaxies where we use strong-line abundance diagnostics, comprising several ground-based spectroscopic surveys targeting Cosmic Noon, a large compilation of archival data observed with JWST, and local galaxies from large spectroscopic surveys.
Across all datasets presented in this work, we applied consistent selection criteria based on emission line signal-to-noise ratios. Specifically, we required S/N $\geq$ 5 for the Balmer lines used in the dust correction (primarily and , but also and when available), and S/N $\geq$ 3 for the lines used in the abundance diagnostics, namely [O ii] and [N ii]. For datasets including auroral line detections, the auroral lines were also required to have S/N $\geq$ 3. When the [S ii] doublet was used in the analysis, it was similarly selected with S/N $\geq$ 3. Finally, we adopted standard diagnostics from Kewley et al. (2013) to exclude potential active galactic nuclei (AGN) contamination. The redshift distribution of the high-redshift sample is shown in Figure 1, demonstrating that the bulk of our objects lie between $z\sim 2$ and $z\sim 4$ (median redshift $z=3.0$).
@@ 2.1.1 The MARTA survey
Figure 1: Redshift distribution of the 664 galaxies for which we measure O/H and N/O in the high-$z$ sample. JWST galaxies with auroral-line detections are shown in light-blue; JWST objects without auroral-line detections (strong-line measurements only) are displayed in teal; objects from ground-based surveys (strong-line measurements from MOSDEF and KLEVER) are plotted in green.